gayparilla:

The price p and the demand x for a particular steam iron are related by the equation x = 1000-20p. express the price p in terms of the demand x and find the domain of this function...

Ultrilliam:

@Hero Please answer this when you can :)

Hero:

"Express the Price of p in terms of x" In other words, solve for p: To solve for p, perform the following steps: 1. Add 20p to both sides 2. Subtract x from both sides 3. Divide both sides by 20 Upon doing so, you should have: \(p = 50 - \dfrac{x}{20}\)

Hero:

The domain in this case will be \(x \ge 0\)

Hero:

Upon further observation, it seems the domain is \(0 \le x \le 1000\). It seems they don't expect the demand to be any higher than 1000

Hero:

Also, if demand is counted in terms of number of units, then the domain would be the set of whole numbers up to 1000